OpenStudy (anonymous):

What is the slope of the line whose equation is 2x – 4y = 10?

OpenStudy (anonymous):

put it in slope-intercept form. (y=mx+b) m is the slope. 2x-4y=10 -4y=-2x+10 y= 1/2x- 5/2 m=1/2 therefore the slope is 1/2

OpenStudy (anonymous):

thank you. u get a medal